OpenStudy (anonymous):

What is the volume of a cone that has a radius of 6 inches and a height of 8 inches?

jimthompson5910 (jim_thompson5910):

Use the formula V = (pi*r^2*h)/3

jimthompson5910 (jim_thompson5910):

In this case r = 6 h = 8

OpenStudy (anonymous):

Okay my answer is 301.44. Is that right?

OpenStudy (anonymous):

what is the volume of a sphere?

jimthompson5910 (jim_thompson5910):

Volume of Sphere V = (4/3)*pi*r^3

jimthompson5910 (jim_thompson5910):

I'm getting 301.44 as well

OpenStudy (unklerhaukus):

jim_thompson5910's formula of the Volume of a cone comes from the combination of the formula a pyramid \(V_\text{pyramid}=\frac13bh\) ( where b is the base and h is the height ) with the formula of a circle \(A_\text{circle}=b=\pi r^2\)

OpenStudy (anonymous):

what is 678.24 / 4/3??

jimthompson5910 (jim_thompson5910):

you're multiplying that with 4/3

OpenStudy (anonymous):

oh! so what would it be? cuz idk how to do that on my calculator

jimthompson5910 (jim_thompson5910):

(4/3)*678.24

jimthompson5910 (jim_thompson5910):

type it in exactly as you see it

OpenStudy (anonymous):

I do, but its not the right answer

jimthompson5910 (jim_thompson5910):

what answer did you get

OpenStudy (anonymous):

something in the 900's?

jimthompson5910 (jim_thompson5910):

904.32 ?

OpenStudy (anonymous):

ya

jimthompson5910 (jim_thompson5910):

this is a sphere with a radius of 6 right? or something else?

OpenStudy (anonymous):

yes a radius of 6, and I have to find the volume

jimthompson5910 (jim_thompson5910):

how do they want you to round? and what do they want you use for pi?

OpenStudy (anonymous):

to the nearest whole number, & 3.14

jimthompson5910 (jim_thompson5910):

ok so 904.32 would round to 904

OpenStudy (anonymous):

Yeah its still saying im wrong...

jimthompson5910 (jim_thompson5910):

hmm

jimthompson5910 (jim_thompson5910):

you'll have to bring this up with your teacher, 904 is the correct answer (according to what they want)

OpenStudy (anonymous):

What is the surface area of a square prism?

jimthompson5910 (jim_thompson5910):

SA = 2*(LW + LH + WH) where L = length W = width H = height

OpenStudy (anonymous):

Okay to double check, whats the surface area for a cylinder?

jimthompson5910 (jim_thompson5910):

SA = 2*pi*r^2 + 2*pi*r*h

OpenStudy (anonymous):

Actually the surface area of a trapezodial prism?

jimthompson5910 (jim_thompson5910):

For any prism, you basically find the areas of the bases, then you add on the lateral surface area (which is basically the sum of the lateral faces)

OpenStudy (anonymous):

Yeah that makes no sense to me

jimthompson5910 (jim_thompson5910):

try looking up a visual example then